### v1

@@ Appendix A Detailed Notes on Individual Clusters
(1) NGC 288: This cluster is classified as G1 based on the presence of tail-like structures extending beyond $r_{\rm t}^{\mathrm{dyn}}$. This result is consistent with previous detections of tidal features in deep DECam imaging from the Dark Energy Survey (2018ApJ...862..114S) and Gaia EDR3 data (2021ApJ...914..123I), and also agrees with the classification of 2020A&A...637L...2P.
@@
(3) Whiting 1: This cluster is likely an accreted member of the Sagittarius dwarf galaxy (2010ApJ...718.1128L; 2019A&A...630L...4M). 2017MNRAS.467L..91C, using MegaCam imaging from both CFHT and Magellan, and 2022ApJ...930...23N, based on DECam photometry from the DECaLS survey, both reported structures extending beyond the tidal radius through matched-filter analysis. In contrast, 2018ApJ...860...66M (also using CFHT/Magellan MegaCam data) and 2022MNRAS.513.3136Z (using deep DECam imaging) found no significant extra-tidal features. Our matched-filter S/N map reveals a tail-like feature extending beyond $r_{\rm t}^{\mathrm{dyn}}$ with an elongated morphology, consistent with the former results. This discrepancy may stem from the higher sensitivity of the matched-filter technique to low-surface-brightness structures. Based on these characteristics, we classify it as a G1 cluster.
@@
(4) NGC 1261: Some indications of tail-like features beyond $r_{\rm t}^{\mathrm{dyn}}$ are present in our matched-filter map, which are further confirmed by the angular density distribution (tail-like features from the lower left to the upper right). This is broadly consistent with previous studies. Based on deep DECam photometry, 2018MNRAS.473.2881K reported a symmetric power-law envelope, and thus PCB20 classified the cluster as G2. However, deeper DES photometry by 2018ApJ...862..114S and Gaia EDR3 data from 2021ApJ...914..123I revealed more extended extra-tidal structures. More recently, 2025A&A...693A..69A used data from the Southern Stellar Stream Spectroscopic Survey (S5) and Gaia DR3, identifying a large number of high-probability extra-tidal member stars beyond $r_{\rm J}$ of NGC 1261. Considering these results along with our own, we classify NGC 1261 as a G1 cluster.
@@
(5) NGC 1851: Previous studies have reported that this cluster hosts an extended envelope reaching roughly $2^{\circ}$ from its center (2009AJ....138.1570O; 2014MNRAS.445.2971C; 2018MNRAS.474..683C; 2018MNRAS.473.2881K). An analysis based on Gaia DR2 (2020MNRAS.495.2222S) did not reveal additional large-scale features, whereas Gaia EDR3 data (2021ApJ...914..123I) suggest the presence of diffuse tail-like structures extending nearly $10^{\circ}$ across the sky. In our matched-filter S/N map, we detect features extending beyond $r_{\rm t}^{\mathrm{dyn}}$, but their appearance is more consistent with a broad, diffuse envelope rather than narrow tidal tails. Given the observed morphology and the depth of our data, we classify this cluster as G2.
@@
(6) NGC 1904: 2018MNRAS.474..683C and 2022MNRAS.513.3136Z reported the detection of a symmetric power-law envelope around this cluster based on deep DECam photometry. However, our matched-filter S/N map reveals clear tidal tails extending well beyond $r_{\rm t}^{\mathrm{dyn}}$. These features are elongated along the north-south direction and are consistent with the structures identified by 2018ApJ...862..114S using DES deep photometry. More recently, 2025A&A...693A..69A also used data from the Southern Stellar Stream Spectroscopic Survey (S5) and Gaia DR3 to identify high-probability extra-tidal member stars beyond $r_{\rm J}$ of NGC 1904. We therefore classify NGC 1904 as a G1 cluster.
@@
(7) NGC 4147: According to the study by 2022A&A...665A...8K, this cluster may be associated with the Sagittarius stream, and 11 extra-tidal member stars were identified, 6 of which lie beyond its Jacobi radius. Earlier, 2010A&A...522A..71J identified tidal tails around NGC 4147 using SDSS photometry, revealing a roughly north-south oriented two-arm morphology, although their data did not extend far beyond the main-sequence turn-off. More recently, 2024AJ....168..237Z, based on DESI photometric data, detected multiple-arm substructures outside the tidal radius. Similarly, our matched-filter results confirm the presence of comparable multiple-arm tidal tails. Considering all these findings, we classify NGC 4147 as a G1 cluster.
@@
(8) NGC 4590: Although PCB20 classified this cluster as G1 based on the detection of tidal stream from Gaia DR2 by 2019MNRAS.488.1535P, our matched-filter S/N map does not reveal any prominent tidal tail features. Instead, we identify a symmetric power-law envelope extending beyond $r_{\rm t}^{\mathrm{King}}$ and approaching $r_{\rm t}^{\mathrm{dyn}}$. Our finding is consistent with that of 2020MNRAS.495.2222S, who used Gaia data and detected a diffuse envelope surrounding the cluster. We find that the large differences in the reported tidal structures primarily result from the different methodologies adopted: study 2019MNRAS.488.1535P integrates the orbit and performs an all-sky search for cluster members, a strategy suited to identifying widely dispersed stellar streams, whereas both our analysis and 2020MNRAS.495.2222S focus on tidal features in the immediate vicinity of the cluster. Given this cluster-centric perspective and the consistency between our result and that of 2020MNRAS.495.2222S, we classify the cluster as a G2 cluster.
@@
(9) NGC 5024: Based on orbital calculations, 2021ApJ...909L..26B suggested that NGC 5024 may be associated with both the Sylgr stream (2019ApJ...872..152I) and the Ravi stream (2018ApJ...862..114S). Previous photometric studies have reported no evidence of extra-tidal features around this cluster (2010A&A...522A..71J; 2012MNRAS.419...14C; 2014MNRAS.445.2971C). A recent wide-field photometric study (2025MNRAS.540.2863W) detected a mild stellar overdensity to the southwest of the cluster, spatially aligned with its orbital path and possibly representing weak tidal debris, but extending no farther than the dynamical tidal radius. Our matched-filter S/N map likewise reveals no significant extra-tidal structure, in agreement with these findings. Based on these results, we classify NGC 5024 as a G3 cluster, consistent with the classification adopted by PCB20.
@@
(10) NGC 5053: Early SDSS-based studies (2006ApJ...651L..33L; 2010A&A...522A..71J) reported tail-like structures, while 2010AJ....139..606C, based on wide-field deep photometry, suggested the presence of an envelope-like component surrounding the cluster and proposed the possible existence of a tidal bridge connecting it to the nearby NGC 5024. Subsequently, 2019MNRAS.485.4906D used Gaia DR2 data to suggest that the tidal structure of NGC 5053 may extend beyond its Jacobi radius. A recent wide-field photometric study (2025MNRAS.540.2863W) detected a mild elongation of NGC 5053 along its western axis and an outer stellar overdensity aligned with its orbital path, but reported no clear evidence of a tidal bridge to NGC 5024. In contrast, some deeper photometric studies (2012MNRAS.419...14C; 2014MNRAS.445.2971C) did not detect significant extra-tidal features, likely due to limited spatial coverage. From the morphology of our detected structure, the cluster shows a faint extension toward the southwest that extends beyond $r_{\rm t}^{\mathrm{dyn}}$. In addition, several diffuse, larger-scale features also lie outside $r_{\rm t}^{\mathrm{dyn}}$. Therefore, we classify this cluster as G2, consistent with the classification adopted by PCB20.
@@
(11) NGC 5272: Based on orbital calculations, 2021ApJ...909L..26B suggested that NGC 5272 may be associated with the Svöl stream (2019ApJ...872..152I). Previous studies based on ground-based photometric data and Gaia observations (2010A&A...522A..71J; 2014MNRAS.445.2971C; 2020MNRAS.495.2222S) have reported no evidence of extra-tidal structures around NGC 5272. A recent wide-field photometric analysis (2025MNRAS.540.2863W) similarly found no clear extension beyond the tidal radius, although it noted the presence of faint stellar overdensities on both sides of the cluster along its orbital direction, which were interpreted as possible tidal tails. Our matched-filter S/N map does not reveal any significant extra-tidal structure, broadly consistent with the conclusions of these studies. Based on these results, we classify NGC 5272 as a G3 cluster, in agreement with the classification of PCB20.
@@
(12) NGC 5466: This is a well-known globular cluster with prominent tidal tails, which have been extensively reported in previous studies based on wide-field photometry and stellar density analyses (e.g., 2006ApJ...639L..17G; 2010AJ....139..606C; 2010A&A...522A..71J). Our matched-filter S/N map clearly reveals symmetric tidal features extending beyond $r_{\rm t}^{\mathrm{dyn}}$ on both sides of the cluster, in good agreement with the morphology reported in earlier studies. Based on this evidence, we classify NGC 5466 as a G1 cluster.
@@
(13) NGC 5634: The study by 2022A&A...665A...8K suggested that this cluster may be associated with the Sagittarius stream, with 10 extra-tidal member stars identified, 5 of which lie beyond its Jacobi radius. Although 2012MNRAS.419...14C; 2014MNRAS.445.2971C provided deep photometric data for NGC 5634, the limited spatial coverage of their observations restricted the evaluation of potential extra-tidal features. More recently, 2025ApJ...988...39W investigated the tidal structure of NGC 5634 using DESI data and found no significant extra-tidal features. Similarly, 2025AJ....170..294C reported no prominent tidal structures based on DELVE DR2 data. Based on the current observational evidence, we classify NGC 5634 as a G3 cluster.
@@
(14) NGC 5897: Studies 2020MNRAS.495.2222S and 2025AJ....170..294C both report extra-tidal features extending beyond the Jacobi radius, though with different morphologies: 2020MNRAS.495.2222S identifies two roughly symmetric extra-tidal features, whereas 2025AJ....170..294C finds only one prominent extension. In our analysis, we likewise detect extra-tidal structures well outside the Jacobi radius. The features we identify are bilateral, and even more pronounced than those reported in 2020MNRAS.495.2222S. Moreover, in the LSST 10-year deep data predictions presented in 2025AJ....170..294C, the two structures appear even more clearly. Based on the structural characteristics revealed in this work, together with the cluster classification criteria adopted in this paper, we conclude that the cluster is more consistent with a G1 type system.
@@
(15) NGC 5904: Although previous studies (e.g., 2010A&A...522A..71J; 2019ApJ...884..174G; 2020MNRAS.495.2222S; 2021ApJ...914..123I) have reported the presence of extended tidal tails around NGC 5904 based on SDSS and Gaia data, our matched-filter S/N map constructed from deep photometric data (reaching $r$-band depth of 23 mag) does not reveal any significant extra-tidal structures. The stellar density distribution appears nearly symmetric and is clearly confined within $r_{\rm t}^{\mathrm{dyn}}$, with no pronounced elongation along any specific direction. A recent wide-field photometric study (2025MNRAS.540.2863W), however, reported strong peripheral deformation, with tidal arm-like features extending north-south and symmetric overdensities along the east-west direction, and suggested that these structures are spatially connected to the cluster’s known rotational axis. In our deeper data, we do not detect such features, indicating that they may possess extremely low surface brightness or be highly diffuse, thus falling below the detection threshold of our matched-filter analysis. Based on our results, we tentatively classify NGC 5904 as a G3 cluster.
@@
(16) NGC 6205: Both 2010A&A...522A..71J, based on SDSS data analysis, and 2020MNRAS.495.2222S, using Gaia DR2 data, did not find any evidence of tidal tails or extra-tidal extensions around NGC 6205. Our matched-filter S/N map similarly shows no significant signs of tidal structures, in agreement with these previous studies. Therefore, we retain the classification of NGC 6205 as a G3 cluster, consistent with the classification by PCB20.
@@
(17) NGC 6229: 2018ApJ...860...66M, based on deep MegaCam photometry, did not identify any significant extra-tidal features around NGC 6229. 2019MNRAS.485.4906D suggested that the cluster might exhibit some degree of extended structure within the theoretical Jacobi radius, based on the fitted density profile. Our matched-filter S/N map also reveals some extended features that clearly extend beyond $r_{\rm t}^{\mathrm{King}}$, though they remain mostly within $r_{\rm t}^{\mathrm{dyn}}$. These features may indicate the presence of marginal outer structures that have previously gone undetected. Based on these results, we conservatively classify NGC 6229 as a G2 cluster.
@@
(18) NGC 6341: 2020MNRAS.495.2222S and 2021ApJ...914..123I identified a long tidal tail associated with NGC 6341 using Gaia data, and this feature was further confirmed by 2020ApJ...902...89T through deep CFHT photometry. A recent wide-field photometric study (2025MNRAS.540.2863W) reported an elliptical extension of stellar density contours beyond the tidal radius, as well as a low-confidence extension aligned with the cluster’s Galactocentric motion vector exhibiting a three-branched morphology, though its physical reality could not be firmly established due to its very low surface brightness. Our matched-filter S/N map also reveals tidal stretching, but the detected features remain confined within $r_{\rm t}^{\mathrm{dyn}}$ and do not extend to larger distances. This discrepancy may be related to the photometric depth limits of the DR10 dataset used in our analysis, as well as differences in the methodological implementations adopted across studies. Based on our observational results, we classify NGC 6341 as a G2 cluster.
@@
(19) Terzan 7: This cluster has long been recognized as a member cluster of the Sagittarius dwarf galaxy (2010ApJ...718.1128L; 2019A&A...630L...4M). According to 2022A&A...665A...8K, a total of 491 extra-tidal member stars were identified in this cluster, 484 of which lie beyond its Jacobi radius. Our matched-filter S/N map reveals clear extended structures on both sides of the cluster, extending beyond $r_{\rm t}^{\mathrm{dyn}}$. The morphology of these features is consistent with typical tidal tails. Based on this evidence, we classify Terzan 7 as a G1 cluster.
@@
(20) Arp 2: This cluster has also long been recognized as a member cluster of the Sagittarius dwarf galaxy (2010ApJ...718.1128L; 2019A&A...630L...4M). In the study of 2022A&A...665A...8K, 315 extra-tidal member stars were identified, including 271 located beyond the cluster’s Jacobi radius. Our matched-filter S/N map reveals a marginal and asymmetric stellar extension, primarily toward the southwest of the cluster, extending beyond $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify Arp 2 as a G1 cluster.
@@
(21) Terzan 8: This cluster is also a member of the Sagittarius dwarf galaxy (2010ApJ...718.1128L; 2019A&A...630L...4M). In the study by 2022A&A...665A...8K, 110 extra-tidal member stars were identified in this cluster, 80 of which lie beyond its Jacobi radius. Our matched-filter S/N map reveals stellar extensions on both the northern and southern sides of the cluster, which extend beyond $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify Terzan 8 as a G1 cluster.
@@
(22) NGC 6864: 2022MNRAS.513.3136Z reported no significant extra-tidal structure around this cluster based on deep DECam data, while 2022MNRAS.509.3709P detected extended features beyond the tidal radius. In our results, we similarly find no clear evidence of tidal tails; however, a envelope-like structure is present beyond $r_{\rm t}^{\mathrm{King}}$ but still confined within $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify NGC 6864 as a G2 cluster.
@@
(23) NGC 6981: Both 2021A&A...646A.176P and 2022MNRAS.513.3136Z suggest that NGC 6981 may possess a low-luminosity envelope. In our matched-filter S/N map, we do not detect tidal tail features, but we do identify an envelope-like structure that extends beyond $r_{\rm t}^{\mathrm{dyn}}$, consistent with the results of these studies. Based on these results, we classify NGC 6981 as a G2 cluster.
@@
(24) NGC 7089: 2010A&A...522A..71J, based on SDSS data, did not find any evidence for extra-tidal structures around NGC 7089. However, 2016MNRAS.461.3639K identified a prominent power-law envelope based on deep DECam imaging, and 2021ApJ...914..123I detected extended tidal tails using Gaia EDR3 data. Our matched-filter S/N map does not reveal obvious tidal tail features, but a possible envelope-like structure is present. Our detection is in better agreement with the results from deeper DECam imaging, and we classify NGC 7089 as a G2 cluster..
@@
(25) NGC 7099: 2020MNRAS.495.2222S reported the presence of prominent tidal tails based on Gaia DR2 data, whereas 2020A&A...643A..15P were unable to reproduce these features using deep DECam photometry, although their analysis revealed a short tidal tail and some scattered stellar debris. In our matched-filter S/N map, we do not detect any clear evidence of extended tidal tails, which may be due to the depth limitations of the DR10 photometric data. However, we do observe hints of peripheral elongation in the outer regions. Based on the current observational results, we classify NGC 7099 as a G3 cluster.

### v2

@@ Appendix A Detailed notes on individual clusters
(1) NGC 288: This cluster is classified as G1 based on the presence of tail-like structures extending beyond $r_{\rm t}^{\mathrm{dyn}}$. This result is consistent with previous detections of tidal features in deep DECam imaging from the Dark Energy Survey (Shipp et al. 2018) and Gaia EDR3 data (Ibata et al. 2021), and also agrees with the classification of Piatti and Carballo-Bello (2020, hereafter PCB20).
@@
(3) Whiting 1: This cluster is likely an accreted member of the Sagittarius dwarf galaxy (Law and Majewski 2010; Massari et al. 2019). Carballo-Bello et al. (2017), using MegaCam imaging from both CFHT and Magellan, and Nie et al. (2022), based on DECam photometry from the DECaLS survey, both reported structures extending beyond the tidal radius through matched-filter analysis. In contrast, Muñoz et al. (2018b) (also using CFHT/Magellan MegaCam data) and Zhang et al. (2022) (using deep DECam imaging) found no significant extratidal features. Our matched-filter S/N map reveals a tail-like feature extending beyond $r_{\rm t}^{\mathrm{dyn}}$ with an elongated morphology, consistent with the former results. This discrepancy may stem from the higher sensitivity of the matched-filter technique to low-surface-brightness structures. Based on these characteristics, we classify it as a G1 cluster.
@@
(4) NGC 1261: Some indications of tail-like features beyond $r_{\rm t}^{\mathrm{dyn}}$ are present in our matched-filter map, which are further confirmed by the angular density distribution (tail-like features from the lower left to the upper right). This is broadly consistent with previous studies. Based on deep DECam photometry, Kuzma et al. (2018) reported a symmetric power-law envelope, and thus PCB20 classified the cluster as G2. However, deeper DES photometry by Shipp et al. (2018) and Gaia EDR3 data from Ibata et al. (2021) revealed more extended extratidal structures. More recently, Awad et al. (2025) used data from the Southern Stellar Stream Spectroscopic Survey (S5) and Gaia DR3, identifying a large number of high-probability extratidal member stars beyond $r_{\rm J}$ of NGC 1261. Considering these results along with our own, we classify NGC 1261 as a G1 cluster.
@@
(5) NGC 1851: Previous studies have reported that this cluster hosts an extended envelope reaching roughly $2^{\circ}$ from its center (Olszewski et al. 2009; Carballo-Bello et al. 2014, 2018; Kuzma et al. 2018). An analysis based on Gaia DR2 (Sollima 2020) did not reveal additional large-scale features, whereas Gaia EDR3 data (Ibata et al. 2021) suggest the presence of diffuse tail-like structures extending nearly $10^{\circ}$ across the sky. In our matched-filter S/N map, we detect features extending beyond $r_{\rm t}^{\mathrm{dyn}}$, but their appearance is more consistent with a broad, diffuse envelope rather than narrow tidal tails. Given the observed morphology and the depth of our data, we classify this cluster as G2.
@@
(6) NGC 1904: Carballo-Bello et al. (2018) and Zhang et al. (2022) reported the detection of a symmetric power-law envelope around this cluster based on deep DECam photometry. However, our matched-filter S/N map reveals clear tidal tails extending well beyond $r_{\rm t}^{\mathrm{dyn}}$. These features are elongated along the north-south direction and are consistent with the structures identified by Shipp et al. (2018) using DES deep photometry. More recently, Awad et al. (2025) also used data from the Southern Stellar Stream Spectroscopic Survey (S5) and Gaia DR3 to identify high-probability extratidal member stars beyond $r_{\rm J}$ of NGC 1904. We therefore classify NGC 1904 as a G1 cluster.
@@
(7) NGC 4147: According to the study by Kundu et al. (2022), this cluster may be associated with the Sagittarius stream, and 11 extratidal member stars were identified, 6 of which lie beyond its Jacobi radius. Earlier, Jordi and Grebel (2010) identified tidal tails around NGC 4147 using SDSS photometry, revealing a roughly north-south oriented two-arm morphology, although their data did not extend far beyond the main-sequence turn-off. More recently, Zhang et al. (2024), based on DESI photometric data, detected multiple-arm substructures outside the tidal radius. Similarly, our matched-filter results confirm the presence of comparable multiple-arm tidal tails. Considering all these findings, we classify NGC 4147 as a G1 cluster.
@@
(8) NGC 4590: Although PCB20 classified this cluster as G1 based on the detection of tidal stream from Gaia DR2 by Palau and Miralda-Escudé (2019), our matched-filter S/N map does not reveal any prominent tidal tail features. Instead, we identify a symmetric power-law envelope extending beyond $r_{\rm t}^{\mathrm{King}}$ and approaching $r_{\rm t}^{\mathrm{dyn}}$. Our finding is consistent with that of Sollima (2020), who used Gaia data and detected a diffuse envelope surrounding the cluster. We find that the large differences in the reported tidal structures primarily result from the different methodologies adopted: study Palau and Miralda-Escudé (2019) integrates the orbit and performs an all-sky search for cluster members, a strategy suited to identifying widely dispersed stellar streams, whereas both our analysis and Sollima (2020) focus on tidal features in the immediate vicinity of the cluster. Given this cluster-centric perspective and the consistency between our result and that of Sollima (2020), we classify the cluster as a G2 cluster.
@@
(9) NGC 5024: Based on orbital calculations, Bonaca et al. (2021) suggested that NGC 5024 may be associated with both the Sylgr stream (Ibata et al. 2019) and the Ravi stream (Shipp et al. 2018). Previous photometric studies have reported no evidence of extratidal features around this cluster (Jordi and Grebel 2010; Carballo-Bello et al. 2012, 2014). A recent wide-field photometric study (Wan et al. 2025) detected a mild stellar overdensity to the southwest of the cluster, spatially aligned with its orbital path and possibly representing weak tidal debris, but extending no farther than the dynamical tidal radius. Our matched-filter S/N map likewise reveals no significant extratidal structure, in agreement with these findings. Based on these results, we classify NGC 5024 as a G3 cluster, consistent with the classification adopted by PCB20.
@@
(10) NGC 5053: Early SDSS-based studies (Lauchner et al. 2006; Jordi and Grebel 2010) reported tail-like structures, while Chun et al. (2010), based on wide-field deep photometry, suggested the presence of an envelope-like component surrounding the cluster and proposed the possible existence of a tidal bridge connecting it to the nearby NGC 5024. Subsequently, de Boer et al. (2019) used Gaia DR2 data to suggest that the tidal structure of NGC 5053 may extend beyond its Jacobi radius. A recent wide-field photometric study (Wan et al. 2025) detected a mild elongation of NGC 5053 along its western axis and an outer stellar overdensity aligned with its orbital path, but reported no clear evidence of a tidal bridge to NGC 5024. In contrast, some deeper photometric studies (Carballo-Bello et al. 2012, 2014) did not detect significant extratidal features, likely due to limited spatial coverage. From the morphology of our detected structure, the cluster shows a faint extension toward the southwest that extends beyond $r_{\rm t}^{\mathrm{dyn}}$. In addition, several diffuse, larger-scale features also lie outside $r_{\rm t}^{\mathrm{dyn}}$. Therefore, we classify this cluster as G2, consistent with the classification adopted by PCB20.
@@
(11) NGC 5272: Based on orbital calculations, Bonaca et al. (2021) suggested that NGC 5272 may be associated with the Svöl stream (Ibata et al. 2019). Previous studies based on ground-based photometric data and Gaia observations (Jordi and Grebel 2010; Carballo-Bello et al. 2014; Sollima 2020) have reported no evidence of extratidal structures around NGC 5272. A recent wide-field photometric analysis (Wan et al. 2025) similarly found no clear extension beyond the tidal radius, although it noted the presence of faint stellar overdensities on both sides of the cluster along its orbital direction, which were interpreted as possible tidal tails. Our matched-filter S/N map does not reveal any significant extratidal structure, broadly consistent with the conclusions of these studies. Based on these results, we classify NGC 5272 as a G3 cluster, in agreement with the classification of PCB20.
@@
(12) NGC 5466: This is a well-known globular cluster with prominent tidal tails, which have been extensively reported in previous studies based on wide-field photometry and stellar density analyses (e.g., Grillmair and Johnson 2006; Chun et al. 2010; Jordi and Grebel 2010). Our matched-filter S/N map clearly reveals symmetric tidal features extending beyond $r_{\rm t}^{\mathrm{dyn}}$ on both sides of the cluster, in good agreement with the morphology reported in earlier studies. Based on this evidence, we classify NGC 5466 as a G1 cluster.
@@
(13) NGC 5634: The study by Kundu et al. (2022) suggested that this cluster may be associated with the Sagittarius stream, with 10 extratidal member stars identified, 5 of which lie beyond its Jacobi radius. Although Carballo-Bello et al. (2012, 2014) provided deep photometric data for NGC 5634, the limited spatial coverage of their observations restricted the evaluation of potential extratidal features. More recently, Wang et al. (2025) investigated the tidal structure of NGC 5634 using DESI data and found no significant extratidal features. Similarly, Chiti et al. (2025) reported no prominent tidal structures based on DELVE DR2 data. Based on the current observational evidence, we classify NGC 5634 as a G3 cluster.
@@
(14) NGC 5897: Studies Sollima (2020) and Chiti et al. (2025) both report extratidal features extending beyond the Jacobi radius, though with different morphologies: Sollima (2020) identifies two roughly symmetric extratidal features, whereas Chiti et al. (2025) finds only one prominent extension. In our analysis, we likewise detect extratidal structures well outside the Jacobi radius. The features we identify are bilateral, and even more pronounced than those reported in Sollima (2020). Moreover, in the LSST 10-year deep data predictions presented in Chiti et al. (2025), the two structures appear even more clearly. Based on the structural characteristics revealed in this work, together with the cluster classification criteria adopted in this paper, we conclude that the cluster is more consistent with a G1 type system.
@@
(15) NGC 5904: Although previous studies (e.g., Jordi and Grebel 2010; Grillmair 2019; Sollima 2020; Ibata et al. 2021) have reported the presence of extended tidal tails around NGC 5904 based on SDSS and Gaia data, our matched-filter S/N map constructed from deep photometric data (reaching $r$-band depth of 23 mag) does not reveal any significant extratidal structures. The stellar density distribution appears nearly symmetric and is clearly confined within $r_{\rm t}^{\mathrm{dyn}}$, with no pronounced elongation along any specific direction. A recent wide-field photometric study (Wan et al. 2025), however, reported strong peripheral deformation, with tidal arm-like features extending north-south and symmetric overdensities along the east-west direction, and suggested that these structures are spatially connected to the cluster’s known rotational axis. In our deeper data, we do not detect such features, indicating that they may possess extremely low surface brightness or be highly diffuse, thus falling below the detection threshold of our matched-filter analysis. Based on our results, we tentatively classify NGC 5904 as a G3 cluster.
@@
(16) NGC 6205: Both Jordi and Grebel (2010), based on SDSS data analysis, and Sollima (2020), using Gaia DR2 data, did not find any evidence of tidal tails or extratidal extensions around NGC 6205. Our matched-filter S/N map similarly shows no significant signs of tidal structures, in agreement with these previous studies. Therefore, we retain the classification of NGC 6205 as a G3 cluster, consistent with the classification by PCB20.
@@
(17) NGC 6229: Muñoz et al. (2018b), based on deep MegaCam photometry, did not identify any significant extratidal features around NGC 6229. de Boer et al. (2019) suggested that the cluster might exhibit some degree of extended structure within the theoretical Jacobi radius, based on the fitted density profile. Our matched-filter S/N map also reveals some extended features that clearly extend beyond $r_{\rm t}^{\mathrm{King}}$, though they remain mostly within $r_{\rm t}^{\mathrm{dyn}}$. These features may indicate the presence of marginal outer structures that have previously gone undetected. Based on these results, we conservatively classify NGC 6229 as a G2 cluster.
@@
(18) NGC 6341: Sollima (2020) and Ibata et al. (2021) identified a long tidal tail associated with NGC 6341 using Gaia data, and this feature was further confirmed by Thomas et al. (2020) through deep CFHT photometry. A recent wide-field photometric study (Wan et al. 2025) reported an elliptical extension of stellar density contours beyond the tidal radius, as well as a low-confidence extension aligned with the cluster’s Galactocentric motion vector exhibiting a three-branched morphology, though its physical reality could not be firmly established due to its very low surface brightness. Our matched-filter S/N map also reveals tidal stretching, but the detected features remain confined within $r_{\rm t}^{\mathrm{dyn}}$ and do not extend to larger distances. This discrepancy may be related to the photometric depth limits of the DR10 dataset used in our analysis, as well as differences in the methodological implementations adopted across studies. Based on our observational results, we classify NGC 6341 as a G2 cluster.
@@
(19) Terzan 7: This cluster has long been recognized as a member cluster of the Sagittarius dwarf galaxy (Law and Majewski 2010; Massari et al. 2019). According to Kundu et al. (2022), a total of 491 extratidal member stars were identified in this cluster, 484 of which lie beyond its Jacobi radius. Our matched-filter S/N map reveals clear extended structures on both sides of the cluster, extending beyond $r_{\rm t}^{\mathrm{dyn}}$. The morphology of these features is consistent with typical tidal tails. Based on this evidence, we classify Terzan 7 as a G1 cluster.
@@
(20) Arp 2: This cluster has also long been recognized as a member cluster of the Sagittarius dwarf galaxy (Law and Majewski 2010; Massari et al. 2019). In the study of Kundu et al. (2022), 315 extratidal member stars were identified, including 271 located beyond the cluster’s Jacobi radius. Our matched-filter S/N map reveals a marginal and asymmetric stellar extension, primarily toward the southwest of the cluster, extending beyond $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify Arp 2 as a G1 cluster.
@@
(21) Terzan 8: This cluster is also a member of the Sagittarius dwarf galaxy (Law and Majewski 2010; Massari et al. 2019). In the study by Kundu et al. (2022), 110 extratidal member stars were identified in this cluster, 80 of which lie beyond its Jacobi radius. Our matched-filter S/N map reveals stellar extensions on both the northern and southern sides of the cluster, which extend beyond $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify Terzan 8 as a G1 cluster.
@@
(22) NGC 6864: Zhang et al. (2022) reported no significant extratidal structure around this cluster based on deep DECam data, while Piatti (2022) detected extended features beyond the tidal radius. In our results, we similarly find no clear evidence of tidal tails; however, a envelope-like structure is present beyond $r_{\rm t}^{\mathrm{King}}$ but still confined within $r_{\rm t}^{\mathrm{dyn}}$. Based on this evidence, we classify NGC 6864 as a G2 cluster.
@@
(23) NGC 6981: Both Piatti et al. (2021) and Zhang et al. (2022) suggest that NGC 6981 may possess a low-luminosity envelope. In our matched-filter S/N map, we do not detect tidal tail features, but we do identify an envelope-like structure that extends beyond $r_{\rm t}^{\mathrm{dyn}}$, consistent with the results of these studies. Based on these results, we classify NGC 6981 as a G2 cluster.
@@
(24) NGC 7089: Jordi and Grebel (2010), based on SDSS data, did not find any evidence of extratidal structures around NGC 7089. However, Kuzma et al. (2016) identified a prominent power-law envelope based on deep DECam imaging, and Ibata et al. (2021) detected extended tidal tails using Gaia EDR3 data. Our matched-filter S/N map does not reveal obvious tidal tail features, but a possible envelope-like structure is present. Our detection is in better agreement with the results from deeper DECam imaging, and we classify NGC 7089 as a G2 cluster..
@@
(25) NGC 7099: Sollima (2020) reported the presence of prominent tidal tails based on Gaia DR2 data, whereas Piatti et al. (2020) were unable to reproduce these features using deep DECam photometry, although their analysis revealed a short tidal tail and some scattered stellar debris. In our matched-filter S/N map, we do not detect any clear evidence of extended tidal tails, which may be due to the depth limitations of the DR10 photometric data. However, we do observe hints of peripheral elongation in the outer regions. Based on the current observational results, we classify NGC 7099 as a G3 cluster.